# Invisible Labor in Open Source Software Ecosystems

### v1

@@ Abstract.
Invisible labor is work that is not fully visible, not appropriately compensated, or both. In open source software (OSS) ecosystems, essential tasks that do not involve code (like content moderation) often become invisible to the detriment of individuals and organizations. However, invisible labor is so difficult to measure that we do not know how much of OSS activities are invisible. Our study addresses this challenge, demonstrating that roughly half of OSS work is invisible. We do this by developing a survey technique with cognitive anchoring that measures OSS developer self-assessments of labor visibility and attribution. Survey respondents ($n=142$) reported that their work is more likely to be nonvisible or partially visible (i.e. visible to at most 1 other person) than fully visible (i.e. visible to 2 or more people). Furthermore, cognitively anchoring participants to the idea of high work visibility increased perceptions of labor visibility and decreased visibility importance compared to anchoring to low work visibility. This suggests that advertising OSS activities as “open” may not make labor visible to most people, but rather lead contributors to overestimate labor visibility. We therefore add to a growing body of evidence that designing systems that recognize all kinds of labor as legitimate contributions is likely to improve fairness in software development while providing greater transparency into work designs that help organizations and communities achieve their goals.
@@
ccs: Software and its engineering Open source modelccs: Software and its engineering Maintaining softwareccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Computer supported cooperative workccs: Social and professional topics Employment issuesccs: Social and professional topics Computing organizationsccs: Social and professional topics Project management techniquesccs: Social and professional topics Software maintenance
@@ 1. Introduction
The processes of creating and maintaining open source software (OSS) involve many diverse tasks beyond just contributing code (Giuri et al., 2010; Eghbal, 2016; Casari et al., 2021; Trinkenreich et al., 2020; Geiger et al., 2021). Conference organizing, community management, finance management, interaction design, data entry, and documenting have all become essential to OSS production and repair (Gousios et al., 2008; Eghbal, 2016, 2020; D’Ignazio and Klein, 2020; Ramin et al., 2020; Young et al., 2021; Trinkenreich et al., 2020). Despite growing interest in supporting noncode contributions (e.g. National Science Foundation, 2023), much of this labor remains undocumented, unattributed to individuals, and “practically invisible” to outsiders and newcomers (Young et al., 2021; Geiger et al., 2021). This often makes noncode contributions forms of invisible labor, undercompensated or uncompensated activities that are “rendered invisible…because they take place out of sight, or because they lack physical form altogether” (D’Ignazio and Klein, 2020).
@@
Labor invisibility creates significant challenges for individuals and organizations involved in OSS. For individuals, lower task visibility makes it difficult to demonstrate the value of their contributions (Suchman, 1995; D’Ignazio and Klein, 2020). In turn, these individuals receive less compensation, whether in terms of future opportunities (Riehle, 2007; Babcock et al., 2017), pay (Toxtli et al., 2021), or social status (Stewart, 2005). And so these individuals frequently burn out and stop performing crucial activities (Eghbal, 2020) that a growing number of scientists, organizations, and governments depend on (Eghbal, 2016). Furthermore, these individuals are often members of minoritized groups (Frluckaj et al., 2022), thereby reinforcing demographic representation disparities (Babcock et al., 2017; D’Ignazio and Klein, 2020).
@@
For organizations and governments, labor invisibility limits work transparency and therefore task measurement, coordination, and control (Suchman, 1995). This limits organizations’ knowledge into whom they fund in open source, what those individuals do, and how those activities shape valued outcomes (Casari et al., 2021). And although invisibility has some benefits for workers in these organizing contexts (like anonymity), computing researchers have long argued that invisibility prevents equity more than creates it (Suchman, 1995; Star and Strauss, 1999; D’Ignazio and Klein, 2020). Indeed, the reliance of open source ecosystems on volunteerism necessitates creating communities where volunteers find benefit in performing all kinds of activities (i.e. not just coding) (Fang and Neufeld, 2009; von Krogh et al., 2012; Eghbal, 2020).
@@
How to make OSS labor (more) visible then? We likely need to (a) recognize diverse types of work that individuals perform and (b) measure how much of those activities people perform. On the former, task taxonomies have seen modest success. For example, academics have started adopting contributorship taxonomies like CRediT to increase attribution and transparency (Brand et al., 2015; Larivière et al., 2016; Holcombe, 2019). Likewise, software developers are experimenting with manual and automated attribution systems (e.g. All Contributors and octohatrack, respectively) with some success (Young et al., 2021). Measuring amounts of labor, though, has proven more elusive, particularly in OSS. Many automated measurement techniques (like GitHub’s contribution activity visualizations) only measure code as “contributions,” thereby conferring lower status to noncoding activities. This provides limited, often overly reductive assessments of “how much got done” (D’Ignazio and Klein, 2020; Young et al., 2021) and can be less meaningful to workers than manual attribution (Monroy-Hernández et al., 2011). And while computational solutions hold potential for spaces like crowd work (Toxtli et al., 2021), they seem unlikely to pass muster in open source ecosystems given the variety of computational and noncomputational interaction contexts involved (coding, conferences, direct chats, etc.).
To address this gap, this work measures invisible labor in OSS ecosystems through two contributions. First, we introduce a technique for measuring labor visibility in open source software development. We use a participatory feminist approach (Bardzell, 2010; D’Ignazio and Klein, 2020) — namely, asking workers directly — to understand their expert perceptions of both how often their work is visible to others and how visible it is to others. We do this by asking developers how often their work is visible to 2 or more people (“fully visible”), to 1 other person (“partially visible”), and to no other people (“nonvisible”). Controlling whether participants see these questions starting from high visibility ($\geq 2$, 1, 0) or low visibility (0, 1, $\geq 2$) with survey randomization (Tourangeau et al., 2013) enables us to estimate labor visibility via cognitive anchoring (Kahneman et al., 1982; Tourangeau, 2018).
Second, we estimate invisible labor in OSS ecosystems by measuring labor attribution alongside our visibility measurement technique. We anonymously surveyed 142 open source software developers across varied fields — from scientific computing to app development — about credit they receive for OSS tasks, their satisfaction with that credit, and the visibility of their work. Randomly giving half of the respondents the descending or ascending orders measured the visibility of the labor they perform and how cognitive anchoring shapes their perceptions. Our results show that across fields roughly half of OSS efforts do not receive credit. OSS activities are more likely nonvisible or partially visible (visible to at most 1 other person) than fully visible (to 2 or more people). Furthermore, individuals anchored to full visibility tended to view their own work as more visible and feel more satisfied with the credit they receive than individuals primed with nonvisibility. This suggests that people in ecosystems with norms of full visibility may overestimate the visibility of their work, making attribution, compensation, and governance inherently more difficult.
@@
We begin by defining labor visibility, invisible labor, and their roles in Computer-Supported Cooperative Work (CSCW). We then describe our methodology and results from a global sample of software developers. We close by discussing what our findings mean for individuals and organizations involved in OSS along with design considerations for overcoming these challenges. Fig. 1 overviews our findings.
@@ 2. Background
To measure something that, by implication, one “cannot see,” it helps to understand what it means for work to be “visible” and “invisible” in the context of OSS. In short, labor becomes invisible when (1) labor data is not fully visible because it does not exist, is not shared, or is not accessible, and (2) labor is undercompensated. The following sections address these concepts in turn.
@@ 2.1. Labor Visibility
When we call work “visible,” we typically do not mean that we can physically “see” the work as it happens. Rather, visibility is a metaphor that captures the availability of information about work, approval of that information for sharing, and accessibility of that information (Stohl et al., 2016; Poster et al., 2016). In labor contexts, availability means that information about the tasks being performed has been created and stored somewhere (e.g. via recordkeeping). Then, those records need to be shared, whether compelled by the law, social norms, or personal conviction, with someone other than the creator. Finally, labor information becomes accessible when an individual knows both where to find the shared information and has sufficient skill to understand what tasks have been performed. Information about labor must meet all three of these conditions — it must exist, be shared, and be accessible — in order for it to be fully visible (Stohl et al., 2016).
@@
2.2. Invisible Labor
Invisible labor has become a popular term to describe marginalized forms of work both inside and outside the home, like childcare (Daniels, 1987), data entry (D’Ignazio and Klein, 2020), and crowd work (Toxtli et al., 2021). In one sense, labor becomes invisible if information about the work being done does not exist, is not shared, cannot be reached, or cannot be understood. Any of these conditions is sufficient to make information about labor, and therefore the doing of labor, invisible (Stohl et al., 2016).
@@ 2.2. Invisible Labor
CSCW has long sought to make labor more visible. Star & Strauss (Strauss, 1985; Star, 1991; Star and Strauss, 1999) coined articulation work [^2] Articulation work is “work that gets things back ‘on track’ in the face of the unexpected [and] is invisible to rationalized models of work,” a classic example of support acts (Star, 1991). in part to raise the visibility of efforts to support distributed work through mechanism design (e.g. new organizing forms, technologies, procedures) (Schmidt and Bannon, 1992; Schmidt and Simone, 1996; Star and Strauss, 1999). More recently, Cherry (Cherry, 2016), Ludwig & colleagues (Ludwig et al., 2018), Gray & Suri (Gray and Suri, 2019), and D’Ignazio & Klein (D’Ignazio and Klein, 2020) revealed other instances of invisibility in computing. Content moderation, proof reading, e-waste processing, and countless digitally-tethered activities struggle with invisibility. In effort to measure invisible labor, Toxtli and colleagues (Toxtli et al., 2021) recently developed a browser plugin to measure invisible labor in crowd work, finding that Amazon Mechanical Turkers spend nearly a third of their time on invisible tasks.
@@
Like with crowd work, OSS development and maintenance activities largely remain invisible (Suchman, 1995; Eghbal, 2016, 2020; Casari et al., 2021), with two consequences. First, labor that deserves compensation is not likely to receive it without greater visibility (Casari et al., 2021). Second, we cannot understand how OSS ecosystems function without greater visibility into the activities of OSS (Suchman, 1995; Casari et al., 2021). Lest we assume this only applies to noncoding activities, consider a coding example too: Projects often combine substantially differentiated coding activities, like bug fixing and feature creation, under the label “code contributions” (Casari et al., 2021). Even activities historically valued as “contributions” are not fully visible for lack of fidelity, thereby making it difficult to assess the health of projects. This further motivates the need for new techniques to measure invisible labor among both code and noncode OSS activities.

### v2

@@ Abstract.
Invisible labor is work that is either not fully visible or not appropriately compensated. In open source software (OSS) ecosystems, essential tasks that do not involve code (like content moderation) often become invisible to the detriment of individuals and organizations. However, invisible labor is sufficiently difficult to measure that we do not know how much of OSS activities are invisible. Our study addresses this challenge, demonstrating that roughly half of OSS work is invisible. We do this by developing a cognitive anchoring survey technique that measures OSS developer self-assessments of labor visibility and attribution. Survey respondents ($n=142$) reported that their work is more likely to be invisible (2 in 3 tasks) than visible, and that half (50.1%) is uncompensated. Priming participants with the idea of visibility caused participants to think their work was more visible, and that visibility was less important, than those primed with invisibility. We also found evidence that tensions between attribution motivations probably increase how common invisible labor is. This suggests that advertising OSS activities as “open” may lead contributors to overestimate how visible their labor actually is. Our findings suggest benefits to working with varied stakeholders to make select, collectively valued activities visible, and increasing compensation in valued forms (like attribution, opportunities, or pay) when possible. This could improve fairness in software development while providing greater transparency into work designs that help organizations and communities achieve their goals.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Software and its engineering Open source modelccs: Software and its engineering Maintaining softwareccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Computer supported cooperative workccs: Social and professional topics Employment issuesccs: Social and professional topics Computing organizationsccs: Social and professional topics Project management techniquesccs: Social and professional topics Software maintenance
@@ 1. Introduction
Creating and maintaining open source software (OSS) involve many diverse tasks beyond contributing code to software repositories (Giuri et al., 2010; Casari et al., 2021; Trinkenreich et al., 2020; Geiger et al., 2021). Tasks like conference organizing, community management, finance management, interaction design, data entry, authorship verification, and documenting have become essential to OSS production, repair, and safeguarding (Gousios et al., 2008; D’Ignazio and Klein, 2020; Ramin et al., 2020; Young et al., 2021; Trinkenreich et al., 2020; Kalgutkar et al., 2019). Acknowledging the breadth and interdependence of these activities, many OSS stakeholders now think about OSS in terms of “ecosystems” (encompassing related software repositories, dependencies, managing organizations, software users, other stakeholders, and their interactions) instead of the less comprehensive idea of “projects” (often limited to code and actions for independent repositories, exclusive of people and interdependencies) (Germonprez et al., 2018; Geiger et al., 2021; Frluckaj et al., 2022; Casari et al., 2023; National Science Foundation, 2023; Jergensen et al., 2011; National Science Foundation, 2024). Shifting our focus from projects to ecosystems — from code to interdependent people and technologies — is important because it reifies a long-held understanding in Computer-Supported Cooperative Work (CSCW) (Suchman, 1995; Star and Strauss, 1999): that much of the labor performed in support of software development is undocumented, unattributed to individuals, and “practically invisible” to outsiders and newcomers (Young et al., 2021; Geiger et al., 2021). In other words, OSS ecosystems involve substantial invisible labor: undercompensated or uncompensated activities that are “rendered invisible…because they take place out of sight, or because they lack physical form altogether” (D’Ignazio and Klein, 2020).
@@
Labor becomes invisible either when (1) labor data is not fully visible (because it does not exist, is not shared, or is not accessible) or when (2) labor is undercompensated (Fig. 1a). While invisibility certainly offers some benefits (like anonymity, autonomy, process flexibility, and reduced coordination), it also creates significant challenges for individuals and organizations involved in OSS (Suchman, 1995; Star and Strauss, 1999). Lower task visibility reduces the extrinsic value of individuals’ contributions (Suchman, 1995; D’Ignazio and Klein, 2020) because individuals receive less compensation, whether in terms of future opportunities (Riehle, 2007; Babcock et al., 2017), pay (Toxtli et al., 2021), or social status (Stewart, 2005). Equally important, it risks delegitimizing less visible tasks and workers, making them more precarious (Star and Strauss, 1999). Consequently, individuals frequently burn out and stop performing crucial activities (Geiger et al., 2021) on which a growing number of scientists, organizations, and governments rely (National Science Foundation, 2024). Furthermore, these individuals are often members of minoritized groups (Star and Strauss, 1999; Frluckaj et al., 2022), which reinforces demographic representation disparities (Babcock et al., 2017; D’Ignazio and Klein, 2020).
@@
For organizations and governments, labor invisibility limits work transparency and therefore task measurement, coordination, and control (Suchman, 1995). This constrains organizations’ knowledge into whom they fund in open source, what those individuals do, and how those activities shape valued outcomes (Casari et al., 2021). For instance, invisibility limits verifiability of authorship (Kalgutkar et al., 2019) among other security challenges (National Science Foundation, 2024), increasingly important concerns in light of the 2024 XZ Utils scandal (Greenberg and Burgess, 2024). In addition, computing researchers have long argued that invisibility prevents equity more than creates it despite the aforementioned benefits (Suchman, 1995; Star and Strauss, 1999; D’Ignazio and Klein, 2020). Indeed, the persisting reliance of open source ecosystems on volunteerism necessitates creating communities where volunteers find benefit in performing all kinds of activities (i.e. not just coding) (Fang and Neufeld, 2009; von Krogh et al., 2012).
@@
Not all work should — or even can — be visible. But given these limitations, some labor certainly should be more visible than it currently is. To make OSS labor (more) visible, we likely need to (a) recognize diverse types of work that individuals perform and (b) measure how much of those activities people perform. On the former, task taxonomies have seen modest success. For example, academics have started adopting contributorship taxonomies like CRediT to increase attribution and transparency (Brand et al., 2015; Larivière et al., 2016; Holcombe, 2019). Likewise, software developers are experimenting with manual and automated attribution systems (e.g. All Contributors and octohatrack, respectively) with some success (Young et al., 2021).
Measuring amounts of invisible labor, though, has proven more elusive, particularly in OSS. Many automated measurement techniques (like GitHub’s contribution activity visualizations) only measure code as “contributions,” thereby conferring lower status to non-coding activities. This provides limited, often overly reductive assessments of “how much got done” (D’Ignazio and Klein, 2020; Young et al., 2021), can be less meaningful to workers than manual attribution (Monroy-Hernández et al., 2011), and confirms concerns of delegitimization. And while computational solutions hold potential for spaces like crowd work (e.g. work on crowdsourcing markets like Amazon Mechanical Turk) (Toxtli et al., 2021), they seem unlikely to pass muster in open source ecosystems given the variety of computational and noncomputational interaction contexts involved (coding, conferences, direct chats, etc.). This leaves a need for research on both the prevalence of invisible labor in OSS ecosystems and factors influencing its prevalence. To address these gaps, this article answers two research questions:
  1. RQ1.
    How common is invisible labor in OSS ecosystems?
  2. RQ2.
    What factors affect how commonly labor invisibility occurs in OSS ecosystems?
We answer these questions by characterizing invisible labor in OSS ecosystems through an explanatory mixed-methods analysis (Creswell and Clark, 2017), with three contributions. First, we introduce a survey technique that estimates labor visibility in OSS ecosystems through cognitive anchoring (Kahneman et al., 1982; Tourangeau, 2018). Next, we demonstrate that roughly half of OSS activities are invisible (Fig. 1b) by measuring OSS labor attribution and visibility across varied fields — from scientific computing to app development. Third, we show that motivations and resource constraints often leave attribution practices at cross-purposes with one another, thereby limiting labor visibility and compensation.
@@
This article begins by clearly defining labor visibility, invisible labor, and their roles in CSCW and OSS ecosystems (Sec. 2). We then describe our methods (Sec. 3), including our participatory feminist approach, survey instrument design, and explanatory mixed-methods analyses. Next, we share results from our quantitative (Sec. 4) and qualitative (Sec. 5) analyses of a global sample of software developers. We close with a discussion (Sec. 6) of what our findings mean for individuals and organizations involved in OSS along with design considerations for overcoming these challenges. Fig. 1 presents an overview of our findings.
@@ 2. Background
To measure something that, by implication, one “cannot see,” it helps to understand what it means for work to be “visible” and “invisible” in the context of OSS. The following sections address these concepts in turn.
@@ 2.1. Labor Visibility
When we call work “visible,” we typically do not mean that we can physically “see” the work as it happens. Instead, we use visibility here as a metaphor that encompasses three qualities: how available information is, how widely shared information is, and how accessible that information is (Stohl et al., 2016; Poster et al., 2016). If a piece of information is fully visible, that information has been recorded, shared widely, and made accessible to many. Likewise, if information in not visible, it must not exist, not have been shared, or not be accessible. [^1] We take the pragmatic view here that if information exists, and is shared and accessible enough (contextually defined), it suffices to call that information “fully visible” for most work contexts. Our definition of visibility is also distinct from verifiability (especially of authorship) which instead focuses on the ability to the author of a specific section of code (Kalgutkar et al., 2019).
In labor contexts, information availability means that data about the tasks being performed has been created and stored somewhere (e.g. via recordkeeping). Then, those records need to be shared, whether compelled by the law, social norms, or personal conviction, with someone other than the creator. Finally, labor information becomes accessible when an individual knows both where to find the shared information and has sufficient skill to understand what tasks have been performed. Information about labor must meet all three of these conditions — it must be recorded, shared, and accessible — in order for it to be fully visible (Stohl et al., 2016).
@@
2.2. Labor Invisibility
Invisible labor has become a popular term for describing marginalized forms of work both inside and outside the home, like childcare (Daniels, 1987), data entry (D’Ignazio and Klein, 2020), and crowd work (Toxtli et al., 2021). In one sense, labor becomes invisible if information about the work being done does not exist, is not shared, cannot be reached, or cannot be understood. Any of these conditions is sufficient to make information about labor, and therefore the doing of labor, invisible (Stohl et al., 2016).
@@ 2.2. Labor Invisibility
CSCW has long taken a dialectical view on labor visibility, seeking both to increase visibility while acknowledging the benefits of invisibility. Star & Strauss (Strauss, 1985; Star, 1991; Star and Strauss, 1999) coined articulation work [^3] Articulation work is “work that gets things back ‘on track’ in the face of the unexpected [and] is invisible to rationalized models of work,” a classic example of support acts (Star, 1991). in part to raise the visibility of efforts to support distributed work through mechanism design (e.g. new organizing forms, technologies, procedures) (Schmidt and Bannon, 1992; Schmidt and Simone, 1996; Star and Strauss, 1999). At the same time, they agree with Suchman’s (Suchman, 1995) arguments that invisibility can also benefit worker autonomy, flexibility, and privacy.
Star & Strauss’ (Star and Strauss, 1999) seminal work discusses two relevant concepts for our discussion. First is the concept of disembedded background work. This describes when workers are visible but the work is not, like how a content moderator may themselves be visible, but the emotion regulation they perform while moderating difficult subject matter may not be. With respect to our definition of visibility, disembedded background work implies that labor data does not exist, is not shared, or is not accessible. Star & Strauss also discuss abstracted and manipulated indicators as another source of invisibility. Here, indicators (like metrics) abstractly describe something that is difficult to measure, meaning the design of these indicators can “hide” work and worker by metaphorically distancing decision-makers from work and worker. For instance, having global pools of users, contributors, and maintainers are often viewed as metrics of success for software projects. Yet, numbers of users or countries (designed measures of success) hide the effort required of many individuals to coordinate global open source organizations as they scale (Geiger et al., 2021). This category (abstracted and manipulated indicators) best matches our concept of partial visibility because labor data — while recorded, shared, and accessible — does not record, share, or make the practical entirety of an individual’s actions accessible.
@@
More recently, Cherry (Cherry, 2016), Ludwig & colleagues (Ludwig et al., 2018), Gray & Suri (Gray and Suri, 2019), and D’Ignazio & Klein (D’Ignazio and Klein, 2020) revealed other instances of invisibility in computing. Content moderation, proof reading, e-waste processing, and countless digitally-tethered activities struggle against the delegitimizing effects of invisibility, suggesting that the detriments of invisibility may outweigh the benefits in many contexts. In effort to measure invisible labor, Toxtli & colleagues (Toxtli et al., 2021) recently developed a browser plugin to measure invisible labor in crowd work, finding that Amazon Mechanical Turkers spend nearly a third of their time on invisible tasks. The pervasive effects of invisibility throughout CSCW raises the question of how invisibility manifests in OSS contexts, as well.
2.3. Invisible Labor in Open Source Software Development
To understand invisible labor in OSS more specifically, we first need to understand a likely source of tension: the varied motivations of those who participate. We can then turn to evidence of invisible labor in OSS ecosystems and factors shaping its prevalence.
2.3.1. Motivations
The most comprehensive review of OSS motivations to date by von Krogh & colleagues (von Krogh et al., 2012) sifts OSS motivations into extrinsic, intrinsic, and internalized extrinsic drivers of participation. Extrinsic motivations are those performed to attain external goods, like pay or career opportunities. Intrinsic motivations — actions performed for their inherent value — encompass ideological, altruistic, kinship (including community-identification), and enjoyment rationales. Participating for reputation, reciprocity, learning, and own-use value are extrinsic motives that participants internalize and perceive as “self-regulating behavior rather than external impositions” (von Krogh et al., 2012, p.653). [^4] For detailed definitions of these motivations, see Carrots and Rainbows: Motivation and social practice in open source software development (von Krogh et al., 2012). We briefly reiterate relevant definitions here for the intrinsic motivations of ideology (that software and its associated practices should be “free” and “open” to all), kinship (supporting those in a group to which one belongs, like an OSS community), and altruism (selfless concern for the welfare of others, regardless of their group affinity or status). Also relevant are career motivations (achieving objectives of interest to a current or future employer). Participants often pursue multiple motivations at once, much like OSS communities do (von Krogh et al., 2012; Sharma et al., 2022; Smirnova et al., 2022). However, motives are commonly in tension with one another because software, social norms, and reward systems differentially advance heterogeneous motivations.
This complicated interplay suggests that individuals might pursue different labor visibility and compensation systems depending on their motivations, which in turn could affect the goal attainment of other participants. For example, one who is motivated by enjoyment alone might not see any benefit in attribution practices, and might even see them as distracting from activities they enjoy. Consequently, they might deprioritize or ignore efforts to increase visibility or compensation. Someone motivated by kinship might see things differently because they seek evidence of their group status (i.e. that they belong or are important to an OSS organization). To them, attribution might confirm their group status, while compensation may or may not. We could similarly conjecture about how visibility and compensation do or do not advance each motivation. Through this lens we can begin to understand evidence of invisibility in OSS.
2.3.2. Evidence of Invisibility
Research to date suggests that much of OSS development and maintenance activities remain invisible (Suchman, 1995; Casari et al., 2021; Geiger et al., 2021). For example, while being an OSS “maintainer” typically describes having access permissions to change code, it might also describe how the individual performs strategic decision-making, code testing, version control, financial accounting, or emotionally-intensive activities like user support, community moderation, user request triaging, relationship management, and advertising (Geiger et al., 2021). [^5] Many of these activities take place “outside of projects” — the traditional activities of creating and modifying code — but are still embedded within ecosystems. However, most of these activities remain less visible (Crowston et al., 2007; D’Ignazio and Klein, 2020; Trinkenreich et al., 2020; Geiger et al., 2021), receive limited financial compensation (Riehle et al., 2014; Trinkenreich et al., 2020), and involve fluid and often circuitous paths for career advancement (Trinkenreich et al., 2020), thereby matching the definition of invisible labor.
@@ 2.3.2. Evidence of Invisibility
The invisibility of such activities affects both participants and organizations invested in OSS ecosystems. Participants who seek compensation for work (whether extrinsic forms like pay or career opportunities, or internalized forms like increased reputation) are not likely to receive it without greater visibility (Casari et al., 2021). Likewise, organizations cannot understand how OSS ecosystems function (or allocate resources appropriately) without greater visibility into the activities of OSS (Suchman, 1995; Casari et al., 2021). Coding examples abound too: projects often combine substantially differentiated coding activities, like bug fixing and feature creation, under the label “code contributions” (Casari et al., 2021), making granular information opaque behind this catch-all term. So even activities historically valued as “contributions” are not fully visible for lack of fidelity, thereby making it difficult to assess the health of projects.
2.3.3. Factors Shaping Invisibility
Several factors might influence the prevalence of invisibility in OSS ecosystems. First and foremost, conflicting motivations seem a likely source of tension. Individuals may align attribution practices with their own motivations, regardless of how widely held those views are in an ecosystem. Properties of cooperative work technologies might also influence labor invisibility via designed features of software development platforms that either track or ignore certain types of tasks (e.g. articulation work) (c.f. Schmidt and Bannon, 1992; Young et al., 2021). Individuals might modify their goals through experiences with technology (Gibson et al., 2022). For example, if an individual feels they are not receiving fair compensation for their efforts, they might seek to increase the visibility of their work in hopes that greater visibility will yield greater compensation. Finally, community (or ecosystem) governance structures could affect labor invisibility. Governance structures strongly incentivize volunteer participation choices (Shah, 2006) with potential to encourage or discourage recordkeeping, sharing, designing for accessibility, and compensating. This may be particularly apparent when stakeholder groups shape norms around conflicting motives, as seen previously in the tension between democratic and bureaucratic governance norms in the Debian ecosystem (O’Mahony and Ferraro, 2007).
While the literature makes a compelling case for the prevalence and potential causes of invisible labor, funders and researchers still argue for more nuanced, granular, and consistent knowledge of labor activities (Casari et al., 2021; Geiger et al., 2021; Frluckaj et al., 2022). The next section describes how we pursue these objectives through a participatory feminist process and explanatory mixed-methods.